OpenStudy (anonymous):

4x - 8y = 20 x intecept

hero (hero):

Hint: Write the given equation in form y = mx + b

OpenStudy (anonymous):

intercept=5/2

hero (hero):

\[4x -8y = 20\] \[4x -20 = 8y\] \[\frac{4x - 20}{8}=y\] \[\frac{x - 5}{2}=y\]

hero (hero):

The x intercept appears to be x = 2y + 5 x = 2(0) + 5 x = 5

OpenStudy (anonymous):

8y = 4x - 20. 4x = 8y + 20. x = 2y + 5. x-intercept = 2(0) + 5, which is 5.

OpenStudy (anonymous):

very good